Cody has had problems starting his two cars during the winter months. He decided to record the number of times each car starts during one month period this winter. He attempted to start each car every morning for 30 days. He recorded the following information:
-His first car started 20 times
-His second car started 18 times
-Both cars started 40% of the time
What is the probability that on any particular morning during the month:
a) at least one of the cars start?
b) he cannot start either of his two cars?

Answer by stanbon(75887)

-His first car started 20 times
P(1st car starts)=20/30=2/3; P(it doesn't start)=1/3
-----------
-His second car started 18 times
P(2nd car starts)=18/30=3/5; P(it doesn't start)=2/5
-----------
-Both cars started 40% of the time
P(both start)=2/5
--------------------
What is the probability that on any particular morning during the month:
a) at least one of the cars start?
P(at least one starts)=1-P(none starts)=1-(1/3)(2/5)=1-2/15=13/15
-------------------
b) he cannot start either of his two cars?
P(neither starts)= 1/3*2/5 = 2/15
Cheers,
Stan H.